OpenStudy (anonymous):

Find the rectangular coordinate of x=cos^2t and y=sin^2t

OpenStudy (anonymous):

@ganeshie8

OpenStudy (perl):

what happens when you add x + y

OpenStudy (perl):

x + y = cos^2(t) + sin^2(t) = ?

OpenStudy (perl):

hint, use pythagorean trig identity

OpenStudy (perl):

also keep in mind that x and y are both positive, since cos^2 and sin^2 are positive. so the graph will be in quadrant 1

OpenStudy (anonymous):

omg got it now

OpenStudy (anonymous):

@perl what about this one x=sin^2t and y=sin^4t

OpenStudy (anonymous):

@satellite73